OpenStudy (anonymous):

What is 3 3/4*(-1 1/2) ? A.3 3/8 B.-5 1/4 C.-3 1/2 D.-5 5/8

OpenStudy (anonymous):

Well you can either distribute or you can turn them into improper fractions and then multiply. Your choice.

OpenStudy (anonymous):

@Susanna

OpenStudy (anonymous):

Hello? Distribute or improper fractions? Which method would you like to use?

OpenStudy (anonymous):

imp

OpenStudy (anonymous):

per

OpenStudy (anonymous):

improper

OpenStudy (anonymous):

Ok, then do you know how to convert a mixed fraction into an improper one?

OpenStudy (anonymous):

\no

OpenStudy (anonymous):

Well what you want to do is turn the whole number next to the fraction, so that it's denominator is the same as the fraction next to it. In \(\large \sf 3\frac{3}{4}\) the denominator of the fraction is 4, so we want the whole number "3" to have that same denominator of "4". But we have to keep the ratio the same, so what we will do is multiply the numerator by 4 as well. Sooo \[\large \sf \frac{3}{1} \times \frac{4}{4}\]

OpenStudy (mathmale):

So, 3 3/4*(-1 1/2) becomes \[\frac{ 15 }{ 4 }*\frac{ -3 }{ 2 }\]

OpenStudy (mathmale):

Simplify this, please.

OpenStudy (mathmale):

If none of the answer choices match your result, perhaps you need to go back and ensure that y ou've copied down the original problem statement correctly.

OpenStudy (anonymous):

here

OpenStudy (anonymous):

so 3/1 * 4*4

OpenStudy (anonymous):

=3

OpenStudy (anonymous):

Yes, that would be the first step to changing it into an improper fraction. When you get the answer, don't simplify it. Keep it in a fraction form.

OpenStudy (anonymous):

i used alculater it simplified for me sorry

OpenStudy (anonymous):

Well we would have \[\large \sf \frac{3 \times 4}{1 \times 4}=?\]

OpenStudy (anonymous):

12/4

OpenStudy (anonymous):

Correct, and now we want to add \[\large \sf \frac{12}{4} + \frac{3}{4}\]

OpenStudy (anonymous):

15/8

OpenStudy (anonymous):

Close, but when we add fractions, the denominator stays the same. So it would actually just be 15/4 not 15/8.

OpenStudy (anonymous):

So now we got the first fraction converted, so let's move onto the second. We want to have \(\large \sf -\frac{1}{1}\) change so the denominator is 2 instead of 1 so we will multiply it by 2/2 and then add on the 1/2 like we did in the last example. So it would be like this \[\large \sf -\frac{1 \times 2}{1 \times 2} - \frac{1}{2}\]

OpenStudy (mathmale):

Please note that we changed the 3 in 3 3/4 to 12/4 so as to be able to add the fraction 3/4. Now both 12/4 and 3/4 have the SAME denominator, which means that you can add them together. As before, add ONLY the numerators; the denominators remain the same and have the value 4.

OpenStudy (mathmale):

So, 3 3/4 becomes 15/4. We need to multiply this by -1 1/2 (minus one and one half). convert this -1 1/2 to a negative improper fraction, please. Then the result you want is (15/4) * (-negative improper fraction from above).